40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog, Page 7 of 58  Not logged in ELOG logo
ID Date Author Type Category Subject
  2626   Wed Feb 3 10:56:54 2021 shrutiUpdatePSOMAcavity locking

I've temporarily removed the phase modulation EOM out of the path to get more power into the cavity.

I have also changed the temperature of the south laser to 7.278 kOhms.

 

  2625   Wed Feb 3 08:23:12 2021 ChrisComputingDAQcymac to Debian buster

I’ll add a note here for future readers of the log, because previous logs on the subject of cymac1 OS upgrades may have made it appear a little too easy.

If you are reading this because you're having an issue with cymac1 (or another front-end computer), you may be wondering: Should I try upgrading the OS on cymac1 to fix my problem?

The answer to that question is: A resounding no!

An OS upgrade is almost never going to be the solution to any problem you may have with cymac1. In fact, it is almost guaranteed to create many new problems that you will have to solve before cymac1 becomes usable again.

Upgrading the OS on cymac1 should only be undertaken if you know exactly why it is absolutely necessary, AND when you have a lot of free time on your hands to troubleshoot the problems it will inevitably create, AND after you have verified there's a backup of the existing system that you can restore to if needed.

UPDATE: the system is running and acquiring data again (although a DAC status bit remains red). Things that had to be fixed to recover from this upgrade were the following:

  1. update daqdrc config file, environment variables in env/systemd_env, and provide symbolic links to additional config files in /etc/advligorts
  2. get new standalone_edc service running to pull in data from external epics channels
  3. rtcds build and rtcds install all models (some had been built but not installed)
  4. add advligorts user to controls group
  5. add group write permissions to files under /opt/rtcds and /frames
  6. get new local_dc service running, to combine data from all front ends for ingestion by daqd
  7. disable mx_crc_only service that is no longer needed

Also to note, rtcds start was working to get front end codes started, but rtcds stop apparently fails to remove the front end modules sometimes (follow it with rmmod when that happens).

  2624   Tue Feb 2 09:52:19 2021 aaronComputingDAQcymac to Debian buster

enter Tue Feb 2 09:52:17 2021

setting up x1oma model

Last time, I ran into symlinking or kernel issues.

on cymac:

sudo apt-get update

Get an error message: GPG error about out of date signatures in the repo, and/or the necessary public key is not available. The error message points to the apt repo's readme, with instructions to add a line to /etc/apt/sources.list.d/lscsoft.list. I uncomment from that file:

deb http://software.ligo.org/lscsoft/debian stretch contrib

but get the same error messages. some of the errors are for stretch-legacy, others for stretch.

I haven't used the realtime system since last year Chris upgraded cymac1's Debian distro from 8 (jessie) to 9 (stretch), and the slow machines to 10 (buster). So, I figure I'll go through the rtcds wiki for setting up cymac with a fine comb, starting by enabling CDSSoftware repos. Buster is now the fully supported Debian release, so to keep all our machines on the same page I'm upgrading cymac1 to Buster. 

install buster

  • Backed up the recommended repos to spirou
  • remove extra cds/lsc software in /etc/apt/sources.list.d (also backed up to spirou), and replace references to stretch in /etc/apt/sources.list to buster
  • apt update to latest version of stretch buster
  • apt-get upgrade to take care of packages that don't require new packages to be removed or installed, then apt full-upgrade to take care of the rest
  • apt install linux-image-4.19-rt-amd64 (hm... was this necessary or desirable? I ended up installing linux-image-amd64 instead, and only later realized that the cymac setup instructions have a unique kernel--see below. Maybe this is related to the kernel errors I was getting earlier?)
  • reboot

Returning to installing cds software:

make a new /etc/apt/sources.list.d/lscsoft.list and cdssoft.list, same as the old but replacing 'stretch' with 'buster'.

apt-get update
apt-get -y --force-yes install lscsoft-archive-keyring
wget -c http://apt.ligo-wa.caltech.edu/debian/pool/buster-unstable/cdssoft-release-buster/cdssoft-release-buster_1.0.5_all.deb
sudo dpkg -i cdssoft-release-buster_1.0.5_all.deb
sudo apt update
sudo apt upgrade  # ifo set to TST and site set to x1 # this step not necessary
sudo apt install cds-workstation # ERROR: dependencies on cds-crtools, ndscope, and python3-nds2-client can't be resolved. error resolved with correction to buster install steps bolded above
# in /etc/apt/sources.list.d/cdssoft-restricted.list add the line
# deb [signed-by=/usr/share/keyrings/cdssoft-unstable-archive-keyring.gpg] https://$USER:$PASSWORD@apt.ligo-wa.caltech.edu/debian buster-restricted main

setting up cymac

sudo apt install advligorts-cymac # same model directory and user
sudo install linux-image-4.19.0-6-rtcds-amd64-unsigned # ERROR: missing destination file operand, the wiki seems to omit apt
sudo apt install linux-image-4.19.0.6-rtcds-amd64 unsigned

The menu structure for this new rtcds kernel is "Debian GNU/Linux>Advanced options for Debian GNU/Linux>Debian GNU/Linux, with Linux 4.19.0-6-rtcds-amd64", which I copy into /etc/default/grub, followed by

sudo update-grub
reboot

Odd, I'm still seeing the old 4.19.0-14-amd64 when running uname -r... which I uninstalled. Removing it again, repeating update-grub, and reboot did the trick. Installation complete.

The only error I'm still seeing is "Skipping acquire of configured file 'main/binary-i386/Packages' as repository 'https://apt.ligo-wa.caltech.edu/debian buster-restricted InRelease' doesn't support architecture 'i386'". This is odd, because I'm running Debian 10 on an amd64 machine, with an amd64 kernel; I last saw this error when configuring cominaux. I think it's because I 'remove'd instead of 'purge'd the old kernels. After purging the non-rtcds kernels, I reboot and... no change in error message. This is not a problem, there is an i386 architecture on cymac1 for running zurich instrument tools, but this shouldn't affect the amd64 architecture; error only occurs for cdssoft-restricted.

outcome

model frontends aren't running. Lab snacks are missing. Outlook grim.

  • Run /opt/rtcds/txt/x1/scripts/startAll.sh. Script runs without error, but I still can't connect to cymac1 through nds.
  • rtcsd status tells me all model are not loaded, and the overall status is 'degraded'. Some kernel modules are missing.
  • rtcds start --all... no change
  • Jon set up a cymac in QIL recently
  • After running systemctl reset-failed, systemctl reports overall state 'running', but running s/startAll.sh gives warnings about identical process variable names on multiple servers. I think reset-failed just silences warnings with no change in state. 
  • /etc/advligorts/env has the same environment variables as before, so problem isn't there
  • This entry from Chris suggests to me perhaps rebuilding all models will help, since they were originally compiled with the old rts
    • rtcds make --all
      rtcds install --all
      reboot
    • no change in systemctl status

  • Investigating the failed systemctl units, I see that rts-daqd is among the units that failed to activate. That's a critical unit, so I inquire with journalctl _PID=1140 (the PID # of the rts-daqd process is listed with systemctl status rts-daqd). The failure is at
    • set master_config="/etc/advligorts/master" 
    • This makes sense, but is surprising--there is no such file, on cymac1 or in my backup! I hadn't heard of this file in any of the rts documentation. The only mention I can find is in a git issue.

It's nearly time for me to hand the lab over to Shruti, so I will stop for now.

exit Tue Feb 2 17:55:24 2021

  2623   Mon Feb 1 18:25:47 2021 shrutiDailyProgressPSOMAcavity locking

Some Observations:

  • Temperature of laser (in kOhms) North: 10.056, South: 10.097
  • Power (in mW) North: 7.02, South: 7.26
  • Beat frequency: ~20 MHz, -50 dBm
  • After changing south laser temperature to 10.985 kOhms, the beat moved to 40 MHz, 4 dBm (!)

Some information:

  • Ring cavity made of :

Ma R~90%, Mb R>96%, Mc R>96% RoC 1m Coastline Optics CO-SAP-1000-250-1.0MP

Ma -> Mb -> Mc -> Ma

  • Photo-diode: PDA20CS

Updates:

  • Got the beam in the triangular cavity to more or less align, i.e., overlap with itself as checked on one side of the triangle
  • I hooked up the transmission PD in the transmission path of Mc
    • At different temperature settings, blocking the incoming beam (Mb -> Mc) resulted in the power becoming minimum and blocking the beam (Mc -> Ma) resulted in some variable fraction of the the nominal power which I take to be a sign of interference (?)
    • I moved around the temperature using the setpoint knob, but since the movement of the knob was nonlinear which along with a nonlinear change in power and frequency (with temperature) made it hard for me to figure out exactly what was happening in the limited time I spent on this so far

Next steps:

  • Going to explore the laser lock on the Moku:Lab for a while, after which I'll decide whether or not to continue with that or move to mixers + PDH servo box

 

  2622   Fri Jan 29 09:52:29 2021 karthikNotes PSOMA Temp Controller tuning

[aaron, karthik]

We met around 930, Aaron gave a brief tour of Cryo, QIL, and EE until 10.

[karthik]

Temp Controller starting 10.981 kOhm, Current 5.66 mA ~29.8 MHz-> Drifting high, ~3.6 dBm

  • Temp knob finicky, use Temp Tune input?
  • PID has no recorded values..

10.963 - 88 mhz

10.954 - 114

10.988 - 11.3 -> 36, check VNA - catching sideband -> switch adjust on/off -> 15.61

Run Series 1

Start at 10.981, step down by small random decrements down to 10.95 and then back up again. Record waveforms into .li file

Logging parameters: duration 10s, 15.6kHz, prefix 29_01_KR

PID is default, see pic

Start End File suffix End freq.
10.981 10.978 104334 43
10.978 10.966 104734 80.1
10.966 10.960 104820 98.8
10.960 10.953 104918 123
10.953 10.963 105213 82
10.963 10.976 105314 50.3
10.976 10.995 105353 6.44
10.981 10.984 105539 35.6-26.1
10.984 10.990 105614 7.3

Stopping here, set parameters back to original, before I play with PID etc, will analyze previous data and iterate from there..

 

 

 

 

  2620   Thu Jan 28 18:03:53 2021 shrutiDailyProgressPSOMAcavity locking

Since the south laser path had the EOM for phase modulation and the north laser path had the intensity modulator, I chose the south for locking to the triangular cavity.

The setup was roughly aligned, but I had to make a few changes:

- The optic labeled 'A' in Attachment 1 had a LH mount which made it nearly impossible to see the cavity reflection. I hunted for a RH mount for the same optic without success, while hunting I decided to switch this R>96% optic to a R~90%.

- The optic labeled 'B' seemed to be a polarizing beam splitter. I switched that to a circular shaped non-polarizing beam splitter.

- I also modified the Y translation knobs of both the fiber launches to make the beam height 4-inches

The image in attachment 1 is modified from elog 2609.

Rana entered the lab and saw that that one of the east table lasers were on and turned it off, after a few minutes I noticed that the south path did not have a beam. Since we did not realize that the south laser was in fact replaced with the beam from the east table, Rana replaced the butterfly mount with one from the cabinet; the cables for the TEC and diode current were also switched out. Once we realized that the south laser was not connected to the current setup, I reconnected it (south laser) replacing the blue fiber from the east table. Both lasers seemed to lase just fine after these changes (after much troubleshooting).

There is no thermal paste on the south laser butterfly mount (may be required later if the PID does not function optimally).

Attachment 1: Path.pdf
Path.pdf
  2619   Thu Jan 28 12:42:14 2021 aaronHowToDAQgrabbing nds data

This is a brief tutorial on grabbing data from the cryo lab using nds2. I'm drawing on Duo's Q measurement script (and and the LIGO DASWG manual and other documentation)

Data in the cryo lab is logged on cymac1, which is only accessible to users within the cryo lab (like the workstations gaston and spirou). I think this means we cannot nds2 without first logging on to a cryo lab computer. Ssh-ing into cymac1 will redirect you to gaston, and the password is the same as for logging into the workstation as controls. 

ssh controls@cymac1.ligo.caltech.edu

A simple python script to grab the current lab temperature with nds2 and print to standard output:

import nds2

TEMP_CHANNEL = 'X1:AUX-LAB_TEMP_F'

conn = nds2.connection('cymac1', 8088)
conn.set_parameter('ALLOW_DATA_ON_TAPE', '1')

for bufs in conn.iterate([TEMP_CHANNEL]):
     temp = bufs[0].data
     print(temp[0])
  2618   Thu Jan 28 11:54:56 2021 PacoNotesEquipment LoanBeam profiler and laptops

[Paco]

Just went into Cryo (~11:54 PT) to drop the profiler and laptops (and chargers). Left them on the left, as you enter the lab, on top of some boxes.

  2617   Wed Jan 27 22:01:16 2021 ranaDailyProgressPSOMAlab improvements

Should we also get some of the 40m kinematic mounts? Maybe get a couple and see if its easier to use.

By 'test equip on the floor' I meant all the stuff under the PSOMA table. That should all go in racks or shelves. Once in awhile the sub-basement labs get a little flooded and we don't want electrical things on the floor for that. Maybe another cabinet on the west wall next to the big grey monster?

  2616   Wed Jan 27 11:58:25 2021 ranaMiscGeneralneeds

You can also use the scripts in https://git.ligo.org/40m/labutils/-/tree/master/moku to directly talk to moku through python or terminal instead of using the Ipad or SD card.

Quote:
 

also, I moved Aaron's large phase tracker data file onto the SD card since the RAM is volatile and was nearly full on the iPad. We should really buy a higher capacity SD card too.

 

  2615   Wed Jan 27 09:11:16 2021 aaronDailyProgressPSOMAlab improvements

enter Wed Jan 27 09:11:13 2021

Facilities mentioned to Paco yesterday that they'd return this morning to complete repairs on the HVAC. I missed him at 9, but around 1pm, Modesto from facilities stopped by to check the air temperature. He measures 63 F at the lab's air intake, and 72 F near the electronics rack. He's checking on our upstream filters to ensure we're getting clean air. Also mentioned that our temperature is centrally controlled, not by the thermostat in the lab -- I was a little confused by this, but sounds like we'll have to contact them to change the temperature. The lab is comfortable for now yes

three corner hat

I'm plotting the three beat note spectra I took Monday them with the help of ctn-scripts. After fixing some improperly set conda permissions on spirou, I didn't untangle its package mess before needing to leave. I've grabbed the files to finish up later.

cable management

I'm putting together a Thorlabs order for cable management and fiber optomechanics -- mounts for our fiber splitters and other pigtailed components.

  • Fiber storage reels (FSR1)
  • fiber component trays, for holding fiber splitters, faraday isolators, etc (BFCT)
  • cable tags (CMS021)
  • Cable tie bases (CMS010) and reusable ties (CMS011)
  • cable straps (for cables that should lay flat) (CS1)

In the meantime, I'm rearranging the electronics rack and cables for increased sanity.

  • Put the L-com BNC panel mount on posts, which may be sturdier than the clamps we had been using
  • Mounting DC power strip to side of electronics rack
    • turn off lasers, then turn off DC power at the Sorensen.
    • Open up power strip to access mounting thru holes on the backside.
    • Mark in pencil the location of 2x holes on the side of the rack, and drill with a #7 gauge bit for #10 rackmount screws (check chart on wall behind EE drill press).
    • Mount the strip's base, then close the power strip as before. Turn on the DC power and lasers.

Fiber launch

The 40m uses these fixed focus collimators (F220APC-1064) in a locking kinematic mount (K6XS) for their fiber launch. Thorlabs also offers similar fixed-focus collimators in 'large beam' (7mm waist for 1550nm) and 'low aberration' varieties. It looks like the triplet lens, low aberration variety uses a slightly different mounting adapter (unthreaded c-clamp, compared with the threaded adapter for F220APC). Thorlabs suggests using epoxy for 'stacked threaded' systems where the internal and external bores both have threads, to avoid inadvertedly loosening one set of threads.

I'm inclined to go with the low aberration version, either T18APC-1550 (3.33 mm waist) or T12APC-1550 (2.27 mm waist). There are also versions of the same with 4.65 and 1.12 mm waist.

 miscellany

  • added a label to the Moku to remind users to plug in ipad after use
  • removing test equipment from floor... we could get something like these electronics cabinets or dolly from McMaster to go under the table. Probably the dolly is more versatile, and we could fit 2 under a table. I've ordered one dolly.

 

exit Wed Jan 27 15:18:12 2021

Attachment 1: Screenshot_from_2021-01-27_10-30-03.png
Screenshot_from_2021-01-27_10-30-03.png
  2614   Mon Jan 25 21:59:34 2021 ranaSummaryLasertest laser

I put the new laser we're testing onto the old Q table on the south side. Its not on yet - we need a Windows laptop to use it. I am thinking that we can use the blue fiber that (right now) runs from the cryo cavities setup over to the PSOMA table. If that one is a  FC/APC connector we can just intercompare the lasers via fiber using the 3-way combiner we talked about.

I had some email exchange with facilities. They seem to be trying to debug the situation remotely and the HEPA on the south wall started blowing cold air sometime around 5-6 PM. Let's see how much the temperature changes (I have a StripTool plot running on the Shruthi workstation).

In order to believe our phase noise measurements, it would be good to engineer some wind/acoustic protection for all teh exposed fiber stuff on the north side of the PSOMA table. Anyone have any ideas?

 

  2613   Mon Jan 25 16:36:24 2021 ranaMiscGeneralneeds

we desperately need:

  1. a sticker on the Moku to make sure the ipad is always left chargin when unused
  2. a rational arrangement and handling of the cables, fibers on the table.
  3. remove all test equipment from the floor under the optical table; maybe get another rack?

also, I moved Aaron's large phase tracker data file onto the SD card since the RAM is volatile and was nearly full on the iPad. We should really buy a higher capacity SD card too.

I am re-running the phase tracker (this time set to frequency instead of amplitude) on the same laser setup. I have tuned the beat frequency to ~30 MHz using the "North" laser temperature.

  2612   Mon Jan 25 13:02:46 2021 aaronUpdatePSOMA 

enter Mon Jan 25 13:02:47 2021

cantilever Q materials

I moved the following materials to measure cantilever Qs from cryo lab to QIL. The are located in bins labelled 'cantilever Q' in a cabinet on the S wall of the QIL lab (see photos).

  • cantilevers
  • HeNe laser and power
  • QPD from the GeNS setup, and power
  • What I think are the clamps Zach was using for cantilever Q measurements.

I also identified in QIL the IR labs cryostat that was used for cantilever Qs, and is now set up for PD measurements. I tried lifting out the cold plate, but felt resistance while opening and stopped. Probably the electrical connections are keeping the cold plate in place, and I would like to check with the user of this cryostat before proceeding.

three corner hat

I am recording spectra of the beat note between three lasers: PSOMA north laser, PSOMA south laser, and cryo cavities East laser. I unplugged the E laser from its fiber-to-free launch (after Faraday isolator), and instead send it to the PSOMA table by fiber. Without changing settings on the cryo cavs TEC or laser driver, the E laser is 8.1 mW when it reaches PSOMA table. When measuring the beat note between the cryo cavs E laser and PSOMA north (south) laser, I send the E laser through the south (north) 90-10 pickoff BS in place of the PSOMA laser. Before recording spectra, I tune the current of the PSOMA N and S lasers until all beat notes are < 20 MHz (there is a few MHz drift during the measurements).

Summary of the measurements, all taken at the FC 1611 (AC measurements on the Moku):

  • acquisition speed: 125 kHz
  • phaselock bandwidth: 10 kHz
  • coupling: AC, 50 Ohm
  • Moku input range: 1 Vpp
  • Measurement time: 1 m 40 s
  • freewheeling: yes
Pair of lasers DC intensity mon (mW) beat note power (dBm) beat frequency (MHz)
PSOMA north and south 6.0 3 18
PSOMA north, cryo cavs east 4.3 -1 24
PSOMA south, cryo cavs east 3.5 -2 5

I notice that the FC 1611 has a UPC connector on its face... but the rest of our components have APC connectors. I'll watch out for effects from this, and order an adapting fiber. I also see sidebands at the harmonics of the beat note when the note is > 0 dBm (linear regime of 1611 is only 1mW).

The rate-limiting step is transferring data from Moku RAM to the ipad over the Wifi... should set up the Moku on ethernet, and check out ctn-scripts for automating the process.

exit Mon Jan 25 16:10:59 2021

  2611   Thu Jan 21 15:25:13 2021 aaronDailyProgressLab Work 

Enter Thu Jan 21 15:25:08 2021

resealed the HEPA filter

over the lab's air intake with Loctite silicone sealant. The rubber cement I used last time delaminated, and I'm hoping this sealant contours to the rough wall paint and proves more effective.

Historically, sealing the HEPA filter decreased the lab temperature. Attachment 1 shows the lab temperature for the last 17 days -- as you can begin to see, sealing the HEPA actually increased the lab filter. This is unfortunate, but consistent with the HVAC blowing hot air despite being set to cool.

 

Organizing electronics rack

I relabelled the 'E' and 'W' path to 'North' and 'South'. This should be less confusing, as the lasers' fiber components run parallel from west-to-east (one laser is north of the other). It would be good to continue labeling, including the AC power cables. I started a techmart cart for cable solutions from digikey (+torque wrench, panel mount SMA feedthroughs, to avoid sending RF through BNC, rack mount components).

Bypassed lossy AOM

The South laser is losing >8dB passing through its AOM. I've taken this AOM out of the path, so the South laser now runs from its Faraday isolator directly to its EOM. For symmetry, I've also removed the North laser's EOM, so the north laser passes through an AOM before being sent to a 90-10 beam splitter. With this modification, the power at the laser launch for the north laser is 7.3 mW and the south is 6.2 mW (the AOM is expected to be lossier than the EOM).

For PSOMA we typically think of our signal being an amplitude modulation, while the pump is phase-modulated for locking. With this modification, the north laser is playing the role of 'signal', while the south laser plays the role of 'pump.' Fine for now, but of course eventually we'll want to swap in a higher power laser for the south (pump) path.

I noticed that the latex gloves being used to softly clamp the AOM have begun to stiffen and deteriorate. Touching the glove at all breaks it into small pieces. This is worrying, so I unmounted the AOM and replaced the gloves. I will look into butterfly mounts and soft, non-degradable cable clamps for the future.

miscellany

  • Pulled aside some bad BNC cables for repair (connector loose, no strain relief).

Exit Thu Jan 21 20:20:27 2021

Attachment 1: Screenshot_from_2021-01-21_19-57-47.png
Screenshot_from_2021-01-21_19-57-47.png
  2610   Fri Jan 15 09:54:18 2021 aaronHowToPSOMAmode matching

[aaron, rana]

aaron enter about Fri Jan 15 09:54:39 2021. rana about

Aaron wiped down keyboard, doorknob, light switch, other high contact surfaces. Rana pointed out several improvements we should make to the optics and electronics, and we tried to collimate the beam coming out of the fiberport.

We noticed the fiber enters the port at an angle, and tried to correct for this to no avail. Turns out, the reason for the angle is to compensate for the angle on the end of FC/APC connectors. You can see an example in the diagram for Thorlabs' fixed focus FC/APC collimators (as opposed to the FC/PC collimators, where the end is perpendicular to the beam and the fiber enters at normal incidence).

We also were concerned that we were seeing only 1.3 mW out of the fiberport -- but we were measuring the W laser path, so this measurement is consistent with the power output from that laser. The E path has 3.9 mW from the fiberport, which I think is consistent with expected losses from the 10% pickoff and couplers.

[aaron]

accounting for observed power

I'm tracing our optical path and accounting for losses. We should at some point rename the E and W lasers, since they are no longer pumping parallel cavities. Pump and signal would be appropriate; cardinal directions are ambiguous, since the relative orientation of the beams changes before the MZ.

Which laser Power after diode (mW) Before isolator After isolator Before EOAM After EOAM Before EOPM After EOPM Before 10% pickoff After 10% pickoff After fiberport
East 21.5 21.5 18.3 18.2 8.4 8.3 4.6 4.6 3.9 3.9
West 20.4 20.5 18.4 18.4 2.4 2.6 1.5 1.5 1.2 1.2

The measurements showing increasing power across connectors (eg after diode to before isolator on W path) are correct, and I attribute to different beam diameter launching from fiber vs from connector affecting the power meter's reading. Not including connectors, these are the expected losses of the components, with anomalous losses highlighted:

Part Expected Optical loss (without connectors, dB) Observed on E path (dB) Observed W path (dB)
IO-G-1550-APC (isolator) 0.55 0.70 0.45
MXAN-LN-10 (EOAM) 3.5 3.36 8.85
MPX-LN-0.1 (EOPM) 2.7 2.56 2.39
PN1550R2A1 (pickoff) 0.8 0.72 0.97

Obviously we should be getting a lot more power from the W path -- and it's encouraging that the laser may not be at fault! Both lasers are near their nominal output power, and in particular we probably could get no more than another 0.1 mW from the E path (assuming we could recover the extra 0.15 dB at the isolator).

I turn off the W laser to inspect the fiber tips with magnifying glasses (from EE), but nothing visibly dirty. Here is an in-depth document describing inspecting and cleaning fibers. Dry wiping the EOAM's pigtails did not change transmitted power. Perhaps we can try again with a fiber microscope for a better view. The west EOAM itself has one bent pin, and is also suspect.

aligning fiberport

Afterwards, I followed Thorlabs' steps to pre-align the W beam fiberport. The E beam fiberport appears collimated to the extent I can tell by eye, but I'd like to use the beam profiler to make fine adjustments.

quick spectrum

Since the fiber PD is set up, I measured the beat note's frequency noise with Moku's phasemeter. Saved the binary and will plot when I'm with my laptop.

exit Fri Jan 15 18:16:38 2021

  2609   Thu Jan 14 17:16:35 2021 aaronDailyProgressPSOMAfree space setup

enter about Thu Jan 14 17:17:03 2021

Finished laying out the MZ, one cavity, and PDs. Need a couple more beam dumps, and should get 3/4'' posts with height appropriate for DC monitor PDs and PBS mount (as in D0901749).

Cleaned up to more easily work with two people tomorrow.

exit Thu Jan 14 19:25:15 2021

Attachment 1: 2CCC0C42-017F-45B5-A4DD-E451953DE4E2.jpeg
2CCC0C42-017F-45B5-A4DD-E451953DE4E2.jpeg
Attachment 2: DF68EBCD-37F3-4CA2-B6C0-6E96146C49C9.jpeg
DF68EBCD-37F3-4CA2-B6C0-6E96146C49C9.jpeg
Attachment 3: 517E4CCF-D12F-4C87-888F-516E21D28049.jpeg
517E4CCF-D12F-4C87-888F-516E21D28049.jpeg
  2608   Wed Jan 13 11:07:24 2021 aaronDailyProgressPSOMAfiber + free 3 corner hat

Picked up Thorlabs and Newport packages from Downs, enter lab Wed Jan 13 11:07:36 2021

The items that arrived today are:

  • FC 1611 photoreceiver, plus 2 power cables
  • cell phone lens kit
  • FC mating sleeve brackets (double only)
  • FiberPort free space couplers, plus mount
  • LB1 beam blocks
  • 90-10 (x2), 50-50 (x1), and 99-1 (x1) fiber beamsplitters
  • rectangular optics mounting bases
  • Remaining ultima mirror mounts (left handed)

Attachment 1 is a week long hour-trend of the temperature at the cymac electronics rack.

I disassembled most of the free space PLL setup, and replaced with fiber components (2x 90-10 BS, 1x 50-50 BS, leading to 1611 and one open ended fiber). On the 90% paths of the first BS, I'm coupling the fibers to free space with the fiber ports.

I also moved the mach zehnder and cavity over, and swapped most of the optic posts-and-forks to post-and-base. Still fiddling with positions / spacings that allow adjustment.

Exit around Wed Jan 13 19:16:39 2021

Attachment 1: labTemp.png
labTemp.png
  2607   Wed Jan 6 10:47:18 2021 aaronDailyProgressPSOMADC to PSOMA rack, realtime model

enter Wed Jan 6 10:47:10 2021 

Making a new cable to run from the Sorensens' rack over to PSOMA rack

  • Grabbed from EE: an appropriately chonky 3-wire cable (AWG 12, same as the old cable), the 18V 'ligo style' connectors (not needed), and some lugs. Confirm the cable is long enough
  • open up the power strip, take a photo to make sure I get the wire mapping right.
  • Remove the old (short) cable, crimp the lugs (and insulator jacket) onto the new cable and install it on the power strip.
  • Run the cable along the cable tray to the Sorensen. Turn off the power supply, then attach the lugs on that end with a 10-32 screw/washer/nut
  • Turn on the supplies, confirm +- 18 V at the power strip. All good, returned parts to EE.

PSOMA realitime model

I figured I'd initialize our realtime model for PSOMA. First, copied over x1cry.mdl (cryo cavities model) into x1oma.mdl. When I build this model, I get the following error

Linux source does not match currently running kernel.

The RCG expects the linux source to be at /usr/src/linux.  For modern
distros this should be a symlink to the source installed as part of
the kernel header package for the running kernel.  For this system:

/usr/src/linux -> /usr/src/linux-headers-3.2.88-rtcds

Please create/update this link, or set the RCG_IGNORE_KERNEL_RELEASE
environment variable to bypass this check

 I haven't seen this before, do I need to update something? I haven't made recent changes to cymac, so I would be surprised if it's running a newer version of linux / rtcds than the existing source (and only one version of rtcds exists in /usr/src anyway). Indeed, when I ask for the currently running kernel (uname -r), it is '3.2.88-rtcds'; and the link at /usr/src/linux points to /usr/src/linux-headers-3.2.88-rtcds. I would expect that the source "linux-headers-3.2.88-rtcds" matches the kernel '3.2.88-rtcds'. I'm setting the ignore flag for now, but would like to know what this means.

Setting ignore to '1' or 'True' does not avoid the error message. I don't see mention of this message on the 40m or other elogs, so I'm a bit stuck. Since the realtime model isn't critical right now anyway, that's all for now.

exit Wed Jan 6 16:27:10 2021

  2606   Tue Jan 5 16:43:44 2021 aaronLab InfrastructureTemperature Sensingfresh temperature sensor

Here is the 10 min trend for the past day. The temperature fluctuates by up to ~0.3 F over short times, and drifted by only ~0.5 F on average throughout the day. The lab is running consistently hot.

Attachment 1: Screen_Shot_2021-01-05_at_16.46.14.png
Screen_Shot_2021-01-05_at_16.46.14.png
  2605   Mon Jan 4 10:46:01 2021 aaronLab InfrastructureTemperature Sensingfresh temperature sensor

Entered Mon Jan 4 10:45:56 2021 .

Temperature in the lab is 89 F at the sensor by the electronics rack. I had been planning to put our AD590s on a board today, unnecessary equipment is already off but I also turned off cominaux and the Sorensen DC supplies. There is hot air coming from HVAC.

I spent some time gathering components and checking out AD's notes on the sensor. I'm using the same components as Andrew/Johannes' old circuit, and putting it on one of Anchal's "generic op amp / diff amp" boards (D1900129-v2). After checking out the old circuit, I realized we still have it in the cryo lab -- it's just missing the AD590. I put the AD590 on a breadboard, made a few connectors, and confirmed that I get a sensible reading (83 K near the center of the room, a couple degrees hotter in/by the electronics rack where the sensor is located).

I noticed that the AC connection on one of our Tenma power supplies is loose (wobbling the cable switches the supply off/on). It's a dual supply, but since I wouldn't want to use it elsewhere I'm powering the AD590 with it (doesn't need to move, out of the way above the rack).

Here's the temperature (second trend) over 20 min, will post the overnight minute trends tomorrow.

Picked up the mess, changed the dust mats, exit

Attachment 1: Screenshot_from_2021-01-04_16-13-53.png
Screenshot_from_2021-01-04_16-13-53.png
  2604   Wed Dec 23 11:13:57 2020 aaronThings to BuyPurchasesoptomechanics arrived

Entered lab around Wed Dec 23 11:14:29 2020 to bring in optomechanics from Newport, step stool from McMaster, and a few other items for around the lab.

  2603   Thu Dec 17 10:23:08 2020 aaronNotesPSOMAlab entry

enter Thu Dec 17 10:24:44 2020

Just deskwork today.

exit Thu Dec 17 16:54:59 2020

  2602   Mon Dec 14 10:32:51 2020 aaronUpdate  

Enter Mon Dec 14 10:32:49 2020

Met with Shruti and finalized the fiber components we want to order.

Alignment procedure has apertures on the E beam path, one before the mixing BS and one after.

  1. Use apertures to determine a path for the E beam
  2. Place the PD such that signal from the E beam is maximized
  3. Align W path to the PD using the penultimate aperture. Because the penultimate aperture and PD position are shared with the E beam, the beams should be copropagating.
  4. Add short focal length lens before the PD.
  5. adjust alignment to maximize beat note.
    1. maximize DC level for E and W with the other path blocked.
    2. maximize beat
    3. repeat

W path laser is down to 25 uW at the PD. Also noticed that the y-alignment depends sensitively on the position of the final lens. The PD height is not the same as the height of the apertures, or the lens is off-center.

Exit Mon Dec 14 15:36:59 2020

  2601   Thu Dec 10 10:05:03 2020 aaronNoise HuntingLaserpll

enter Thu Dec 10 10:04:54 2020

Quote:

Doesn't the phase meter just read out the noise even with no locking? I thought that was going to be the magic.

For locking, the mixer readout is in units of phase and the laser current modulation gives a proportional frequency modulation with no frequency wiggles until > 1 MHz. So it should phase lock with no integrator, but I'm not sure if the free running noise will drive it out of the phase lock or not. I wonder if its possible to use the phase meter as an error signal. It would be much easier to lock frequency instead of phase via a mixer.

Hm, hadn't tried the phasemeter application. I'll check it out now... if I understand your second comment, you're saying because

f = \frac{d\phi}{dt} \\ \implies \int f dt = \Delta \phi

an error signal proportional to phase is already integrating the frequency error. Makes sense, but does 'easier to lock frequency instead of phase via a mixer' follow, or is that unrelated? 

The Moku phasemeter does produce a nice power spectrum. Here it is up to 200 Hz, I'm working with Anchal's ctn-scripts and pymoku to get the higher frequencies. 

Still odd that the beat amplitude is so small. Let's check:

quantity Power of E beam @ 1611 (power meter) Power of W laser @ 1611 (power meter) 1611 DC voltage from E beam 1611 DC voltage from W beam DC voltage gain responsivity @ 1550 (approx) Expected DC voltage due to E beam Expected DC voltage due to W beam
value 307.8 uW 65.7 uW  -1.55 V  -400 mV 10 V/mA 1 A/W -3.08 V -660 mV

Looks like neither beam is producing the expected photocurrent, but because the error is not the same factor for both beams I suspect alignment / beam size. I'm aligning with some apertures to avoid smearing the beam on lenses. Aligning each beam led to more power, but my procedure doesn't simultaneously align both beams. 

exit Thu Dec 10 15:11:30 2020

Attachment 1: IMG_0016.PNG
IMG_0016.PNG
  2600   Sun Dec 6 19:22:10 2020 ranaNoise HuntingLaserpll

Doesn't the phase meter just read out the noise even with no locking? I thought that was going to be the magic.

For locking, the mixer readout is in units of phase and the laser current modulation gives a proportional frequency modulation with no frequency wiggles until > 1 MHz. So it should phase lock with no integrator, but I'm not sure if the free running noise will drive it out of the phase lock or not. I wonder if its possible to use the phase meter as an error signal. It would be much easier to lock frequency instead of phase via a mixer.

 

  2599   Fri Dec 4 10:36:18 2020 aaronNoise HuntingLaserpll

Entered lab, then grabbed a spool of cable from EE, started elog Fri Dec 4 10:37:52 2020

thought about filters. The narrowest line I managed (yes really) is in the attached screenshot. I amplified +40 dB with Agilent 8447A before the splitter.

exit Fri Dec 4 16:14:19 2020

Attachment 1: IMG_0015.PNG
IMG_0015.PNG
  2598   Thu Dec 3 11:57:52 2020 ranaNoise HuntingLaserW laser oscillating

Definitely always use 50 Ohm for all inputs and output with RF frequencies. The Moku should be able to drive the current modulation input of the diode driver to directly phase lock based on the 1611 signal with no amplifiers. A 1 mW beat should give you ~1 Vpp, which is a few dBm.

  2597   Thu Dec 3 10:18:07 2020 aaronNoise HuntingLaserpll

Entered somewhat before Thu Dec 3 10:18:07 2020

finishing up the PLL. I still need to set an appropriate gain for the LO, but in the meantime I'll try to use the Moku's laser lock feature

Moku

This is pretty straightforward. Moku has an internal oscillator and lets you control the LP (corner frequency) and controller filter (proportional gain, integrator frequency, differentiator frequency, integrator saturation level, differentiator saturation level). I'm driving the E laser HF and LF inputs from the Moku outputs. Quickly acquire a lock and play around with filter settings for a while.

exit Thu Dec 3 12:30:47 2020

  2596   Wed Dec 2 13:37:55 2020 aaronNoise HuntingLaserW laser oscillating

Enter Wed Dec 2 13:37:57 2020

I'd like to know whether the problem is in the W laser driver (S1500207) or the diode (104987). I swap the E laser drive over to the W laser diode, so am driving diode 104987 with driver S1600246. I still don't see the E laser driver oscillating, but the W diode power still reaches only 1.38 mW before falling off (ramping up current at T_set). Just in case, I also drive the W diode using the E driver but the cable from the original W driver (checking that the cable isn't the cause); no oscillation.

Entered EE to obtain grounding strap (Wed Dec 2 14:10:12 2020). On second thought, putting off diagnosing the W laser drive and just continuing with ITC 502. I did get a look inside, and don't actually see evidence that R33 was changed, just a added resistor from R30 to V_ref mentioned on the schematic.

Found this elog helpful for making sure the pinout for ITC 502 is compatible. Oddly, I can't set the laser diode current limit higher than 10mA, though the range of the device is +- 200 mA. I suspect the left 2 digits of the current display are just inoperable, because I can turn the adjustment pot in either direction until the digits stop moving. I set the current limit at 149 mA by counting by 10, and will double check the eventual current from the monitor. the diode definitely is nearing end of life, I'm only getting 1mW at nominal setpoint. I was squeezing a little extra out of the custom laser driver, but probably at the expense of the diode.

With the thorlabs driver, we're down to 1 peak at last!

Looking at the spectrum today, maybe the low frequency cutoff I noticed yesterday is just a frequency offset issue?

switching over to Moku... wow this is a user friendly device.

  • beat note drifts by MHz over minute timescales
  • FWHM ~ 140 kHz
  • 35 dB above noise floor

Phaselock setup

  • Photodiode output to splitter (ZFSC-2-4+)
    • Long cable to 83 dB (!!?) gain amplifier to mixer (ZAY-1B) LO input. Mixer requires +23 dBm LO.
    • short cable to [gain] to mixer (ZAY-1B) RF input

While choosing the gain for producing an LO from the 1611 output, I tried to measure the absolute power with the Moku. When I change the range from 1 Vpp to 10 Vpp, the overall peak height changes by almost 10 dB. I wondered if this were just a clipping issue, so I hooked up the Moku function generator output to the channel 2 input to see if the behavior is consistent. The beat note is 300 mVpp at the laser settings I'm currently using, which should be about 0 dBm dissipated into 1 MOhm. Sourcing 300 mVpp from the Moku into its channel 2 input gives the correct power reading, but the 300 mVpp from the 1611 reads -25 dBm with 10 Vpp range and -34 dBm with 1 Vpp range. What's going on?

  • Probably just outside the linear range of the 1611, which saturates at 1mW input power. I'm seeing some unexpected sidebands (>> 9MHz), so nonlinearity makes sense. We're still well below the 10 mW damage threshold, but should add back the 90-10 pickoff mirrors now that alignment is done.
  • After adding the 90-10 (power) BS, the AC output of the 1611 has ~3 mVpp (amplitude). nice
  • The beat is now -53 dBm. That's still below what I'd expect for a 3 mVpp signal; Moku thinks it's seeing a 1.45 mVpp
    • I want to attribute this to some setting of the oscilloscope or Moku, but it's also not consistent with my expectation. moku shows 6 mVpp when measuring its 3 mVpp output with 1 MOhm input impedance, suggesting there's a factor of 2 I'm missing somewhere; on the other hand, the signal from the 1611 needs a factor of 1/2 to be consistent with the oscilloscope measurement. Awkward.
  • started these measurements at 50 Ohm input impedance for the SA, but when results didn't make sense I repeated at 1 MOhm. Should have stuck with 50.

I would have to string together eg 2x SR560 to get enough gain. I guess that's fine, because the noise floor of the SR560 is O(nV/rtHz)... but it gives me the feeling of doing something wrong, and I'm hungry so will pick it up tomorrow.

exit Wed Dec 2 18:42:15 2020

  2595   Tue Dec 1 10:59:49 2020 aaronNoise HuntingLaserpll

Entered Tue Dec 1 10:59:48 2020

Turning on the lasers in a more controlled way today, trying to reach datasheet nominal setpoint

  1. turn down current driver to 0mA
  2. Turn on TEC (P and D at 0, only I => definitely no oscillating).
  3. Ramp up the current to the datasheet's I_bias, while monitoring the DC current mon and laser power. According to the schematic, the total current mon is 1 V / 10 mA.
    • Zach mentioned some modifications to this board in an elog. Hm, I've been using the schematic's current mon mapping (max 100mA), so maybe I've already damaged one of the diodes and that's why we're having issues. I don't see a record of the second (W) laser drive being modified, so perhaps this is why we are only seeing the 10MHz line on the E laser HF current mon.

I'm getting turned around, so I'll summarize the state of the drivers and lasers. Yellow highlight indicates this is a best guess based on things like dates on the DCC, but I haven't verified by eye (by opening the driver chassis or making a measurement).

Laser Driver serial number Driver part number known Modifications Voltage limit DC current mon Measurements
East S1600246 D1200719-v4 none no 1V / 10mA (this must be wrong) Zach 2017
West S1500267 D1200719-v3 current doubled no 100 V/A Zach 2017

Something seems wrong with the W laser path. At the nominal laser setpoint, the E path puts out a steady 5.33 mW; the W path puts out up to 1.9ish mW, but the power is fluctuating between 1 and 2 mW.

 

Spent some time changing the W/E mixing BS into a michelson BS for the W path (uneven arms). The AS beam from one leg was substantially brighter (by eye and ~10x on the PD) than the other. I confirmed that the mirror is HR for 1550. Probably just clipping, I had the plate BS kinematic mount in the wrong handedness to avoid remounting it; this was misguided anyway, I return it to original state. When I realigned the PLL path (identical to before this Michelson excursion), the forest of modes returned to the gaussian envelope state (not the bessel 2 looking envelope from yesterday). Could this be alignment / path length dependent? I returned the lasers to nominal T and I, and the gaussian envelope remains, so optical path is my best guess.

A little later, I lowered the TEC setpoint for W laser, and the Bessel envelope returned. However, whereas yesterday the 2nd sideband had a maximum now the 1st sideband is maximized.

Another feature that's been puzzling me -- when I sweep the temperature or current monotonically in a direction that moves the beat to 0Hz, the forest moves towards 0 until about 50 MHz. Below 50MHz, the modes are suppressed nearly to the noise floor; I think the carrier is just visible above the floor, but above 50Mhz the carrier is 50 dB above the floor. The cutoff is sharp, and if I continue sweeping temperature or current in the same direction the modes eventually reappear above 50 MHz moving up. My guess is it's another 'feature' of the analog spectrum analyzer that I haven't worked out (maybe secretly normalizing out the 1/f? but it's faster than 1/f rolloff), and that something cuts off low frequency sensitivity. Seeing as I'm well within 200MHz, I'm switching to the moku to check.

While the ipad charges, made this table of the modes I'm seeing at the nominal T_set of 23 C (10.940 kOhm) for W laser, 25 C (10.050 kOhm) for E laser. The marker tells me sideband spacing is 9.6 MHz; the W current drive HF mon has a line at 9.7 MHz, so it does seem these are related. I've attached the oscilloscope trace, where you can see that the W laser drive HF mon (chan 4) has RMS noise at least 100x the noise on E laser HF mon. The oscillation is dominated by the peak at 9.7 MHz, though there are a few others. Maybe the solution is just to swap in another laser driver -- this driver is a modified version an out of date revision of the circuit. Tomorrow I will swap in the combi controller for the W current driver and see if that helps.

E laser current (mA / 2) W laser current (mA / 2) Envelope order carrier frequency (MHz) # of sidebands (100kHz RBW)
77.41 69.50 0 155 3
75.54 69.5 0 177 3
53.6   0 995 3

 

exit Tue Dec 1 18:08:59 2020

  2594   Mon Nov 30 11:28:30 2020 aaronNoise HuntingLaserpll

Entered lab about Mon Nov 30 10:21:35 2020, after taking a COVID test through Caltech's new surveillance testing program.

I'll pick up where Shruti left off on the beat note. The comb of sidebands becomes a single line  remains a comb when the PID is offKoji suggests maybe the (PLL) PID is oscillating at 10Mhz.

  • Turned on both laser drivers, TEC, 15V PD power supply, HP 8560E spectrum analyzer.
    • E laser TEC was set to 8.301 kOhm
    • W laser TEC was set to 8.992 kOhm
  • Located the thermistor resistance on the RIO data sheets
    • Diode #104978 (E path) has thermistor resistance 10.050 kOhm at the nominal setpoint (25 C)
    • Diode #104987 (W path) has thermistor resistance 10.940 kOhm at the nominal setpoint (23 C)
    • Oddly, Dmass' elog on unboxing these lasers lists a different operating temperature than the datasheets
  • Adjusted the PID on both TEC
    • First, turned down I and D to 0. Next, increased P until there was oscillation. Added back a little D and turned down P until no oscillation. Increase I until it moves to the setpoint.
    • Checked 'step response' when the laser is turned on, and adjusted PID until no overshoot or oscillating on the way to setpoint, and equilibrium reached within seconds. This did not affect the sidebands.
    • Adjusting P while moni
    • toring the beat spectrum results in no change in the sidebands, until the temperature starts to oscillate and the beat is lost.
  • Checking out the HF current mon on a 200MHz oscilloscope to see if something's going on there. The HF drive input is open, so there should be no peaks.
    • W laser has a line at 9.75 MHz, 20 dB above the noise floor.
    • E laser has a line at 22.5 MHz, 20 dB above the noise floor (on second look, this peak is not present. Lots of adjustment between the two measurements though, and I'm not sure what did the trick)
  • I'm seeing the forest of sidebands hopping side-to-side (~38 MHz hops). Modehopping? [no, it was a triggering issue] I adjust the TEC setpoint to the nominal values on the datasheet.
    • The hops are always by exactly 4 times the sideband spacing (~40MHz, so the second to the right sideband frequency at +20 MHz coincides with the second to the left sideband frequency at -20 MHz after the hop)
    • Hm... turning off both lasers, I see several noise peaks in the dark spectrum that 'turn on and off' every other refresh of the spectrum video...
      • This is present even when I take single spectra manually (not continuous). Every other spectrum has some noise forest around 20MHz. Is this an artifact of the SA settings? No reason something at MHz should exhibit such repeatable on/off behavior for my random-near-Hz button presses.
        Indeed, when I turn on the 'frequency counter' on the spectrum analyzer, the hopping goes away... why?
      • With frequency counter on, I now see two sets of sideband peaks. What's the spectrum analyzer doing for demodulation, this looks like some artifact. Perhaps some saturated stage of the spectrum analyzer is causing sidebands. Here's a nice technical note from HP on spectrum analyzers.
  • Larger changes of temperature and current moves the forest of peaks out to higher frequency (after being lost for intermediate values of T and I).
  • The forest of modes is still there even with both PID off.

 

exit Mon Nov 30 16:12:04 2020

Attachment 1: IMG_0195.jpg
IMG_0195.jpg
  2592   Mon Nov 30 10:26:17 2020 ranaPhotosSi fabSi cantilever photos

Thanks, the photos are now on the shared drive.

  2590   Sat Nov 28 21:53:12 2020 ranaPhotosSi fabSi cantilever photos

For storing lab photos in W Bridge, you can use our shared google acct instead so that we all have access to it (see chat for secrets)

  2589   Fri Nov 27 19:43:33 2020 KojiUpdateLab Workpsoma locking

Isn't the PID oscilating at 10MHz?

  2588   Fri Nov 27 13:38:18 2020 ShrutiUpdateLab Workpsoma locking

Attachment 1: Video of spectrum analyzer with zoomed out beat after turning off the PID loop of west laser

Attachment 2: Another image of the zoomed in spectrum when the PID is on.

Attachment 1: Beat.mp4
Attachment 2: BeatZoomed.pdf
BeatZoomed.pdf
  2587   Tue Nov 24 15:58:00 2020 aaronNotesEquipment Loanmoku CTN to Cryo

Anchal dropped off the Moku from CTN, along with its USBA->USBC cable, power cable, and ipad.

  2586   Tue Nov 24 13:23:09 2020 aaronPhotosSi fabSi cantilever photos

Entered lab around Tue Nov 24 13:24:57 2020 to finish photographing Zach's cantilevers.

some things about cameras, and in particular the FinePix F300 EXR

  • ISO -- the camera's sensitivity to light. More sensitive means more noise, but also more signal (useful when exposure time must remain short). 
  • aperture -- size of the opening before the lens. A wide aperture yields a shallow depth of field and lets in more light, but can cause blurriness in the foreground and background. Narrow aperture lets in less light and widens the depth of field, but can lead to diffraction effects or not enough exposure time depending on the application.
  • shutter speed -- how long with the shutter remain open? All the usual tradeoffs of integrating.
 
For these shots, I have the camera mounted on a tripod and close enough to the cantilevers that the subject takes up the full field. I've turned off the overhead lights and oriented the bright, fluorescent desk lamp away from the camera and slightly up. I'm reflecting some diffuse light back to the cantilevers with a large kim wipe (and my white face / lab coat). I've set minimum ISO, long exposure time, and large aperture. Since I'm handling nominally clean Si, I'm wearing gloves, lab coat, mask, and hair net. I covered the wall behind the shot with paper towel to provide a dark, uniform surface that isn't visible in reflections off the wafer containers and doesn't backlight the shot. The photos are not taken from above because doing so resulted in reflection from the top cover, which I wanted to keep on to avoid dust.
 
Looks like 5-7 cantilevers could be high Q (no visible contaminants, pitting, cracks, etc). I think there's also a pair of high Q cantilevers in Zach's cantilever cavities cryostat, and maybe one in the cantilever op lev cryostat (QIL). If we measure the Q of the most promising bare cantilevers, we can identify the 3ish best candidates for aSi coating.
 
Photos can be found in this album. I've pulled a representative good-looking cantilever and attach them here, along with the photo booth setup.
 
attachment 1: the photo booth
attachment 2: the lines near the top are reflections from the edge of the container. The long sides are parallel, and can give you a sense of the angle of the photograph. 
 
While doing this, I entered EE to retrieve a spare battery for the camera, and later again to return the camera to its place by the sink.
 
left around Tue Nov 24 17:50:48 2020
Attachment 1: DSCF3580.JPG
DSCF3580.JPG
Attachment 2: DSCF3572.JPG
DSCF3572.JPG
  2585   Tue Nov 24 10:16:04 2020 shrutiUpdateLab Workpsoma locking

Updates

- Found a suitable power cable M-M for the New Focus 0901 power supply on the east table (I did not realize yesterday that these were the same cables). Then I checked the voltage on the pins and they were fine.

- Using the New Focus 1611 (1 GHz PD) powered by the New Focus 0901 +-15 V / 0.3 A max. power supply, I tried finding the beat note. I looked at the RF output on a HP 8560 E spectrum analyzer and the DC output on an oscilloscope.

The DC output ranged from 500 mV to over 1 V as I scanned the temperature of one or both lasers.

- When the east laser temperature read roughly 8.34 kOhm and west was 9.04 kOhm I saw a pattern as in Attachment 2.

Changing the temperature slightly did cause the peaks to shift about, and further when I changed the polarization of the east laser using the HWP the height of the peaks varied. They also disappeared when either of the beams were blocked.

The estimated peak power in the taller peaks is ~0.1 µW from the plot.

- I also tried scanning the temperature of both lasers again to possibly find a single peak. No luck yet.

Today, I didn't check the alignment very carefully and I probably have to tune it further after the changes that Aaron and I made over the past few days.

The next step is to do the phase-locking.

 

 

Attachment 1: BeatSetup.pdf
BeatSetup.pdf
Attachment 2: BeatOrNot.pdf
BeatOrNot.pdf
  2584   Fri Nov 20 19:24:00 2020 aaronElectronicsElectronicspower cables

Entered lab Fri Nov 20 19:24:32 2020, usual sanitation.

[blue "Photodiode Power Supply] Looking for a DB9 to BNC adapter. I found this spider instead -- close enough. Use multimeter to measure 24V between pints 4 an 9... not promising. Confirm power is connected, no signal on the frontpanel BNCs either. Could remove this one and take a look on the benchtop, but above is...

[Newfocus +-15V current-limited power supply] Has 3 pole bananas and a power switch on the front. Found a power cable for the back. There's a bananas to 3-pin LEMO already there. Double check the voltage with a multimeter. Alas the connector doesn't fit the PD, but should be some cables in the EE shop or elsewhere...

Didn't find the right connector in EE. On the 'power cables' rack (NE corner Cryo), there was a M-F connector, but I need M-M. Could cannibalize the 12VDC supply? I think for now +-12V is working, so should look a bit more.

 

I moved the power cables for our preamps for better strain relief (attachment 2 is the before photo). 

I also had left this ND filter sitting on the table (attachment 1). Yikes!

More photos here.

ExitFri Nov 20 20:39:43 2020

Attachment 1: DSCF3541.JPG
DSCF3541.JPG
Attachment 2: DSCF3542.JPG
DSCF3542.JPG
  2583   Thu Nov 19 16:34:01 2020 aaronDailyProgressGeneralcantilever photos, laser intensity noise

Entered lab Thu Nov 19 16:33:57 2020. Usual sanitation, personal reminder to report campus access with Caltech.

Cantilevers

we want to coat some of Zach's cantilevers with a-Si so we can make a cold Q measurement. I've started to take some photos, but have become tired and will finish tomorrow. There are O(5) suitable cantilevers produced in January 2018, but I'll have to dig a bit more (or ask Zach) to determine what's what. We can measure the Q of the most promising few cantilevers to be sure they're acceptable.

I borrowed the digital camera from EE shop, but left its case (which is very dusty).

Laser intensity noise

Sending W path beam to a Newfocus 1811 to measure free running laser intensity noise.

Following Shruti's recent diagram, I moved the Newfocus 1811 into position after OMTL1. I also moved PO1.1 back into the beam path, so I can use it to align into the 1811. Turn on the E laser and TEC, also had to move Ma for alignment. I still don't have a +-15V power supply, will ask around. Turn off the laser and TEC before exit at Thu Nov 19 20:41:22 2020

  2582   Thu Nov 19 11:23:44 2020 shrutiUpdateLab Workpsoma locking

Attachment 1: An updated version of the diagram in elog 2577 where the path lengths to the beat beam-splitter are identical. The fiber launchers and some components have been moved around, but everything after PO1.1 along the beam has been retained as before.

Attachment 2: Retaining the same configuration to the beat BS, the cavity with Mach-Zehnder interferometer has been added. Also the path lengths to the MZ input BS along both laser beam paths have the same length. Except for the ring cavity, the Mach-Zehnder is also balanced.

Attachment 3: Updates pertaining to the current setup

  • Work in progress to achieve the configuration in Attachment 1.
  • I have switched the two PDs so the beat can be measured with the Newfocus 1611 (has a larger BW) and the noise measurement with the east laser can be done with the Newfocus 1811.
  • I removed the 10/90 pick-off circled in green, so that would have to be added to the path to continue the noise measurement.

 

Attachment 1: PLL_FS_sym.pdf
PLL_FS_sym.pdf
Attachment 2: MZ_PLL_FS_sym.pdf
MZ_PLL_FS_sym.pdf
Attachment 3: NewBeat.pdf
NewBeat.pdf
  2581   Wed Nov 18 08:24:02 2020 shrutiUpdatePSOMAtemp control and transverse beam profiles

D [in.]

West laser (X) [um]  West laser (Y) [um] East laser (X) [um] East laser (Y) [um]
2 327.2, 336.5 332.0, 340.9 327.1, 332.8 330.0, 335.2
3 358.2, 269.6 363.4, 374.7    
4 420.4, 439.3 417.8, 439.8    
5 511.7, 549.2 510.2, 549.9 527.6, 550.7 519.6, 540.4
6 642.8, 688.0 630.6, 674.9    
7 766.4, 807.9 754.6, 801.4 778.0, 848.3 716.0, 782.7
8 891.1, 932.6 894.7, 942.1    
9 974.3, 1023.1 943.9, 1000.6    
10 1142.6, 1193.9 1152.7, 1203.5 1092.2, 1159.7 1040.6, 1103.4

 

This is the data using the Data Ray Beam'R2 profiler with the InGaAs window. Attachment 1 contains images of each of those profiles.

D: distance from fiber launcher in inches; The two values in each of the cells are [Clip 13.5%, 4 sigma] respectively, i.e., the method used to calculate the beam widths.

The previous measurement using a razor blade refers to 'sigma' which I believe explains why these values are 4 times larger.

These profiles were taken with temperature stabilized such that the powers were ~1 mW.

East laser set to 8.070 k Ohm, West laser set to 9.065 k Ohm. I don't understand why there is such a difference.

 

Other updates:

I had hooked up the ITC 502 combi controller to the west Rio laser and used only its temperature controller. (I believe both the thermistors that measure the diode temperatures are TH-20k Ohm.)

Both the PID controllers work satisfactorily: the TED 200 C with the east laser stabilizes to within few Ohms of the setpoint thermistor resistance within some seconds;

the ITC 502 stabilizes at a similar rate but at an offset of ~10 Ohms despite the integrator being set to maximum. I fiddled around with the P and I settings a little but realized that this configuration seemed optimal.

 

To measure these profiles at different distances I moved the fiber launcher head and then replaced it back to its original position, roughly.

Attachment 1: beamprofdata.zip
  2580   Tue Nov 17 14:56:06 2020 aaronDailyProgressGenerallab entry

I entered the lab somewhat before Tue Nov 17 14:56:22 2020. Exited Tue Nov 17 16:39:12 2020

Hand sanitizer on entry, also sanitized the bulky green laser goggles before and after my use (forgot contacts). Turned off the laser, sat at desk and considered turning on the laser. Took a break. On my walk I wrote this haiku

Took another break.
Why align these mirrors when
photons have no mass?
 
I'm going to this seminar, so that's all for today. If nothing else, a reminder (even to myself) to always elog, no matter how small. 
  2579   Mon Nov 9 14:40:46 2020 shrutiMiscEquipment LoanStuff for N2 transfer

I've placed the following items outside the cryo lab:

1. Cryo liquid N2 dewar

2. Funnel

3. Tube

4. Two pairs of cryo gloves

Attachment 1: IMG_0334.pdf
IMG_0334.pdf
  2578   Wed Oct 28 18:10:27 2020 aaronNoise HuntingLaser

Lab entry

in: Wed Oct 28 18:11:00 2020
out: Wed Oct 28 21:05:58 2020
 

E laser noise

  • Moved the Newfocus 1611 from W path cryo cantilevers transmission, to just after the PSOMA beam launch.
    • nominal AC gain: 700 V/A
    • nominal DC gain: 10 kV/A
    • output impedance: 50 Ohm
    • Max input power: 1 mW DC (10 mW CW)
  • Measuring the following spectra across the PD band (~30kHz-1GHz) by connecting the 1611 AC output directly to an HP 8560E (50 Ohm input). I'm monitoring DC with TDS3024B oscilloscope (1 MOhm input).
    • dark current
    • free running E path laser intensity noise

Optical layout: beam launch -> lambda/2 -> steering mirrors -> lens 1 -> ND 0.6 -> lens 2 -> PD 1611

There is only a 12 VDC power supply compatible with the 1611 power port, but the PD requires +- 15V. Surely there's one somewhere. Perhaps this is why I observe only -6 V on the DC mon with 1mW input power at 1550nm (checked against the Thorlabs S122C; I expected -10V). Maybe the beam is too large.

  2577   Wed Oct 28 12:42:31 2020 shrutiDailyProgressLab Workpsoma locking

Phase-locking the two lasers:

Updates:

- Although when we talked about adjusting the MZ-phase, we decided that having the phase/path length control with fiber components might be better initially (Refer Attachment 2), for now I began doing everything in free-space.

- Attachment 1 shows the setup as it is now. Previously I'd placed polarizing beam splitters instead of 90/10 beam-splitters because I thought it would be easier to work with, but now changed my mind and decided to stick with what we planned.

Next steps:

(Once the beat is obtained on the spectrum analyzer)

1. Adjust set-point temperatures to adjust beat frequency to the right frequency.
Since ideally we want the two frequencies to be identical, it might be a good idea to add an AOM to one of the paths so that the first order beam is mixed with the other laser and this beat be compared to a stable reference for phase-locking.
But, initially we plan to skip the AOM.
2. Adjust the PID parameters if needed
3. Add electronic components

- Measure the laser frequency noise


29 Oct 20:

I've added Attachment 3 -- which is the current free space version and some PLL electronics. 

- It does not show the Mach-Zehnder part as that will be added only later

- This setup is asymmetric but in a future version we will change that

Attachment 1: Setup2020Oct27.pdf
Setup2020Oct27.pdf
Attachment 2: psoma_PLL.pdf
psoma_PLL.pdf
Attachment 3: psoma_pll_freespace_intermediate.pdf
psoma_pll_freespace_intermediate.pdf
  2576   Tue Oct 27 15:57:27 2020 shrutiDailyProgressLab Workpsoma locking

Today I modified the optical setup with the aim of obtaining the beat between the two diode lasers for phase-locking.

I added pick-off polarizing beamsplitters with HWPs in each path for now (to be able to adjust their power) and mixed them at a 50/50 non-polarizing beam-splitter to eventually reach a Newfocus 1811 low noise PD.

I will add pictures and more details later.

 

  2575   Fri Oct 23 19:06:05 2020 gautamUpdateEquipment LoanTwo resonant RFPDs ---> 40m

I was in the Cryo lab between 1215-1230 this afternoon. I removed two resonant RFPDs from what was Johannes' setup (encircled in Attachment #1). I also brought a SR554 preamplifier to the 40m.

I was wearing the usual PPE (gloves, face mask) while I was in the lab.

Attachment 1: IMG_8903.JPG
IMG_8903.JPG
  2574   Thu Oct 15 22:15:11 2020 Ian MacMillanComputingSimulationModeringer Simulation

This is the spectrum coming off of the sample. there should be a peak at 1038Hz... but there isn't. And what is even weirder is that the spectrum analyzer that is built into Simulink shows a peak where I expect but when I do it here it doesn't show up. 

Update: I think I have found why there is a discrepancy between the two versions of the power spectrum. The spectrum analyzer in the Simulink model requires non-continuous data so you have to use a block to make your data discrete. The sampling rate of that block affects where the peak of the mode is seen. So it seems that that the mode seen in the previous post was just caused by making the data non-continuous.

 

Attachment 1: ModeRingerSpec.pdf
ModeRingerSpec.pdf
ELOG V3.1.3-